OpenStudy (anonymous):

2sin^2x+sinx=1 solve for x

OpenStudy (anonymous):

\[2x^2+x-1=0\\(x+1)(2x-1)=0\\x=-1,x=\frac{1}{2}\] now replace \(x\) by \(\sin(x)\) and solve \[\sin(x)=-1\] and \[\sin(x)=\frac{1}{2}\] for \(x\)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw